# Generated Contents Enrichment

### v2

@@ Abstract
In this paper, we investigate a novel artificial intelligence generation task, termed as generated contents enrichment (GCE). Different from conventional artificial intelligence contents generation task that enriches the given textual description implicitly with limited semantics for generating visually real content, our proposed GCE strives to perform content enrichment explicitly on both the visual and textual domain, from which the enriched contents are visually real, structurally reasonable, and semantically abundant. Towards to solve GCE, we propose a deep end-to-end method that explicitly explores the semantics and inter-semantic relationships during the enrichment. Specifically, we first model the input description as a semantic graph, wherein each node represents an object and each edge corresponds to the inter-object relationship. We then adopt Graph Convolutional Networks on top of the input scene description to predict the enriching objects and their relationships with the input objects. Finally, the enriched description is fed into an image synthesis model to carry out the visual contents generation. Our experiments conducted on the Visual Genome dataset exhibit promising and visually plausible results.
Index Terms:
Text-to-Image Generation, Scene Graph, Graph Neural Networks, Generative Adversarial Networks
I Introduction
@@
Given the rich semantics and complex inter-semantic relationships in the real-world scenes in Fig. LABEL:fig:teaser(f), their available descriptions, shown in Fig. LABEL:fig:teaser(a), are usually inadequately informative, thus leading to the semantic richness gap between the accordingly generated images and the real-world ones, as shown in Fig. LABEL:fig:teaser(d) and (f). However, given the simple descriptions, our human minds can easily hallucinate a semantically abundant image, filling the semantic richness gap by first understanding the given description, then enriching it with its related semantics and inter-semantic relationships, and finally hallucinating the accordingly generated image.
@@
We term this task, barely investigated in conventional research, as generated contents enrichment (GCE). Although the human brain can fulfill this task effortlessly, existing artificial intelligence content generation methods produce results with inadequate semantic richness. In contrast to our approach, these methods overlook explicit content enrichment by neglecting reasoning on relevant semantics and inter-semantic relationships. As shown in Fig. LABEL:fig:teaser(d), the state-of-the-art generation approach [2], without reasoning the related semantics of the given descriptions, generates the images with insufficient semantics compared with the real-world ones in Fig. LABEL:fig:teaser(f). Indeed, all existing content generation methods, to the best of our knowledge, rely on implicit learning to generate visual contents that follow the distribution of visually real images without explicitly considering the semantic richness and structural reasonableness, thus incompetent in tackling GCE.
@@
Although GCE is essential for filling the semantic richness gap between the short description and the expected image, it is challenging to learn to solve it. Firstly, the enriching content should be semantically relevant to the ones in the description and form appropriate inter-semantic relationships. Secondly, the enriched scene graph, composited by the description and its enriched semantics, should be structurally real, which follows the distribution of scene graphs from real-world images.
@@
Towards solving the mentioned challenges and achieving GCE, we propose an intriguing approach that, to some extent, mimics the human reasoning procedure for hallucinating the enriching content. Particularly, our introduced method is designed with an end-to-end trainable architecture consisting of three principal stages. As an initial step, we represent the input description as a scene graph, wherein each node resembles a semantic, i.e., an object, and each edge corresponds to an inter-object relationship. In the first stage, we enrich the scene graph iteratively. In each iteration, we append a single object by modeling it as an unknown node and its related inter-object relationships as unknown edges, which are then predicted employing a Graph Convolutional Network (GCN) by aggregating the information from the known nodes and edges; thus, the enriching contents are semantically relevant with the ones in the description. Then, the enriched scene graph is fed into a pair of Scene Graph Discriminators to guarantee its structural reality. In the second stage, we synthesize the enriched image from the enriched scene description by adopting an image generator [3, 2]. The third stage simultaneously feeds the enriched image and the input description into the Visual Scene Characteriser and Image-Text Aligner, respectively, thus the generated image is visually real and preserves the information of the description.
@@
Therefore, our contribution is a novel end-to-end trainable framework appointed to handle the proposed GCE task as the first attempt, to our best knowledge. This is executed by demonstrating the input scene descriptions as scene graphs and then enriching the semantics by employing GCNs. Afterward, the final enrichments are produced by feeding the enriched scene descriptions into an image generator. In contrast to traditional text-to-image generation approaches, which solely focus on visual authenticity, we enhanced the generation by incorporating explicit visual and textual content enrichments that are semantically and structurally coherent with the provided input description. We assess our method on the real-world images and scene descriptions with rich semantics from the Visual Genome dataset [1], and achieve encouraging quantitative and qualitative outcomes.
II Related Work
II-A Realistic Image Generation
@@
Generative Adversarial Networks (GANs) [4, 5] jointly train a generator and discriminator to distinguish real from fake synthesized images. GANs produce images of high quality that are, in some cases, almost impossible for humans to discriminate from real images [6, 7, 8, 9]. However, their training procedure is not as stable as the other methods, and they may not model all parts of the data distribution [10]. The first attempts for GAN-based text-to-image synthesis [11, 12] are followed by stacking GANs [13] and employing the attention mechanism [14, 15, 16]. Variational Auto Encoder (VAE) [17] consists of an encoder to a latent space and a decoder to return to the image space. VAE is also employed for image synthesis [18, 19], generating high-resolution data but with lower quality than their GAN counterpart. Alternatively, autoregressive approaches [20, 21, 22, 23, 24] synthesize pixels conditioned on an encoded version of the input description that could depend on the previously generated pixels with computationally expensive training and sequential sampling process. Although state-of-the-art diffusion-based [25] models [26, 27, 28, 29] synthesize images of astounding quality, these architectures necessitate elevated computational costs during training and a time-consuming inference. Stable Diffusion [2] is one of the first open-source methods that attracted so much attention among artists as it applies the diffusion process to low-dimensional image embeddings. On the other hand, there has been some research on a hybrid two-stage method like VQ-VAEs [30, 31] and VQGANs [32, 33] which aims to benefit from a mixed version of the mentioned methods.
II-B Graph Neural Networks (GNNs)
Neural networks were first applied to graphs in [34] and further explored in subsequent studies such as [35, 36, 37, 38]. Following improvements in computer vision tasks stemming from the introduction of CNNs, convolution-based models for processing graphs were proposed and divided into two main categories: spectral-based [39, 40, 41, 37, 42] and spatial-based [43, 44, 45, 46]. GNNs may handle many different tasks [47], such as clustering the graph nodes [48], predicting edges between nodes [48], and partitioning the graph [49]. Other GNN applications include but are not limited to text or document classification or capturing the internal structure of text [50, 51], machine translation [52], recommender systems [53, 54], programming [55, 56], predicting social influence [57], preventing adversarial attacks [58], brain networks [59], combinatorial optimization [60], detecting events [61], chemistry-related tasks such as understanding the molecular fingerprints [62], predicting a structure’s properties [46], generating new chemical structures [63, 64], and even predicting the taxi demand [65]. Specifically in Computer Vision, GNNs are employed to tackle human activity recognition [66, 67], processing 3D point clouds [68, 69], interactions of humans and objects [70], image classification with only a few samples [71], question answering [72], visual reasoning [73], and semantic segmentation [74].
II-C Scene Graphs
@@
Scene graphs proposed in [75] and further investigated in [76, 77, 78] are a structured version of a scene representation containing information about objects, their relationships, and often their attributes. Previously, understanding a scene was limited to detecting and recognizing the objects in an image. However, scene graphs aim to obtain a higher level of scene understanding [79] and are employed to improve action recognition [80], image captioning [81], and retrieve images [82]. Also, scene graphs may be predict from images [83, 84, 85, 86, 87, 88] and serve as inputs to generate synthesized images [3, 89, 90, 91, 92].
II-D Semantic Expansion
@@
The task of visual comprehension is addressed by generating an enriched and expressive scene graph as output from a given input image [93, 94]. This is performed by incorporating common-sense knowledge from the input images. Additionally, [95] focuses on the task of image outpainting, where an input image is expanded beyond its original boundaries to output an extrapolated version of the input. Their approach involves constructing and extending a scene graph based on the input image.
In contrast, our work introduces and tackles the task of GCE, aiming to enhance the image synthesis quality. While previous studies have engaged in semantic expansion, their objectives, input-output structures, and evaluation criteria differ significantly from ours. Specifically, the tasks addressed in the literature involve generating scene graphs from images, whereas our GCE approach involves transforming scene descriptions into enriched descriptions and generating their corresponding enriched images.
III Method
@@
Figure -1: High-Level Overview of our end-to-end Generated Contents Enrichment framework during the training phase. In Stage 1, the input scene graph (SG) is fed to the Scene Graph Enricher ${G_{sg}}$ to produce an enriched SG. Besides, a pair of local and global discriminators in the SG Critic $D_{sg}$ are jointly trained to differentiate between original and enriched SGs. These discriminators aid the enrichment process in constructing realistic, structurally coherent, and semantically meaningful scenes. In Stage 2, the Image Synthesizer $G_{im}$ leverages the resultant enriched SG to generate an image. In Stage 3, essential visual and textual scene characteristics are extracted in the Visual Scene Characterizer $S_{cf}$ and the Image-Text Aligner $M_{im\_sg}$. These two components ensure that the enriched image appropriately reflects the original description’s inherent characteristics.
@@
Fig. -1 illustrates a high-level representation of our end-to-end model during the training phase. The model accepts a scene graph (SG) $g$ as input to simultaneously enrich the content in both textual and visual domains. Our proposed approach for GCE consists of three stages performed iteratively.
@@ III-A Scene Graphs
III-B Graph Convolutional Network
Our Graph Convolutional Network (GCN) shares similarities with Convolutional Neural Networks (CNNs) in various aspects. Though GCNs are primarily designed to process graphs as inputs, they have the capacity to produce an output graph with a structure similar to the input. This is analogous to CNNs employing appropriate paddings. Each GCN is composed of multiple graph convolution (GConv) layers utilized to aggregate the data from nodes and edges, enabling the propagation of information throughout the graph. In this framework, the $l^{th}$ GConv layer, denoted as $GConv^{(l)}$ performs the message-passing as:
@@ III-B Graph Convolutional Network
In the context of the $GConv^{(l)}$ input graph, $V_{O}^{(l)}$ represents the feature vectors for all nodes, and $V_{R}^{(l)}$ denotes the feature vectors for all predicates. Additionally, the corresponding output vectors are characterized as $V_{O}^{(l+1)}$ and $V_{R}^{(l+1)}$. To initialize the process, $V_{O}^{(0)}$ and $V_{R}^{(0)}$ serve as the initial vector values for the objects and predicates within the GCN input graph.
Like CNNs, our approach employs weight sharing. This design choice allows the GCN to accommodate graphs with an arbitrary number of vertices and edges. Moreover, parallelization is feasible due to the independence of calculations for each node, enhancing computational efficiency.
The fundamental building block of our GCN is the GConv layer. For every object $o_{i}\in O$ and edge $(o_{i},r,o_{j})\in E$, we concatenate the three corresponding sets of feature vectors, which consist of $(v_{i},v_{r},v_{j})$ representing the subject, predicate, and object vectors, respectively. These concatenated feature sets are then passed through the GConv layer for further processing, initiating with an MLP represented as $f_{g}$:
$\begin{split}(\bar{v}_{i},\bar{v}_{r},\bar{v}_{j})=f_{g}(v_{i},v_{r},v_{j}).\end{split}$ (2)
This is to compute the set of three intermediate subject, predicate, and object vectors $(\bar{v}_{i},\bar{v}_{r},\bar{v}_{j})$. Subsequently, these intermediate vectors are aggregated to form $V_{O}^{(l)}$ and $V_{R}^{(l)}$. Node features $V_{O}^{(l)}$ include the new feature vectors $v^{\prime}_{i}$ for the output nodes and edge features $V_{R}^{(l)}$ consist of the new feature vectors $v^{\prime}_{r}$ for the output predicates:
$v^{\prime}_{r}=f_{r}(v_{r},\bar{v}_{r}),$ (3)
$\begin{split}C_{i}^{s}=\{\bar{v}_{i}|(\bar{v}_{i},\bar{v}_{r},\bar{v}_{j})=f_{g}(v_{i},v_{r},v_{j})\text{ and }(o_{i},r,o_{j})\in E\},\\ C_{i}^{o}=\{\bar{v}_{i}|(\bar{v}_{j},\bar{v}_{r},\bar{v}_{i})=f_{g}(v_{j},v_{r},v_{i})\text{ and }(o_{j},r,o_{i})\in E\},\end{split}$ (4)
$v^{\prime}_{i}=f_{o}(v_{i},\frac{1}{|C_{i}^{s}|+|C_{i}^{o}|}\sum_{\bar{v}_{i}\in C_{i}^{s}\cup C_{i}^{o}}\bar{v}_{i}).$ (5)
@@
For more comprehensive information regarding the structure of our GCN and GConv, including detailed illustrations and further explanations, please refer to Section Method and Fig. 1 in the supplementary material.
III-C Stage 1: Generative Adversarial SG Enrichment
III-C1 SG Enricher.
The Scene Graph Enricher, serving as our SG generator and featured in Fig. -1, encompasses three distinct steps, each tailored to address specific subproblems associated with appending a new object, edges, and predicates. These steps leverage the proposed GCNs along with an Enriching Edge Detector. Here is an overview of the process:

### v3

@@ Abstract
In this paper, we investigate a novel artificial intelligence generation task termed Generated Contents Enrichment (GCE). Conventional AI content generation produces visually realistic content by implicitly enriching the given textual description based on limited semantic descriptions. Unlike this traditional task, our proposed GCE strives to perform content enrichment explicitly in both the visual and textual domains. The goal is to generate content that is visually realistic, structurally coherent, and semantically abundant. To tackle GCE, we propose a deep end-to-end adversarial method that explicitly explores semantics and inter-semantic relationships during the enrichment process. Our approach first models the input description as a scene graph, where nodes represent objects and edges capture inter-object relationships. We then adopt Graph Convolutional Networks on top of the input scene description to predict additional enriching objects and their relationships with the existing ones. Finally, the enriched description is passed to an image synthesis model to generate the corresponding visual content. Experiments conducted on the Visual Genome dataset demonstrate the effectiveness of our method, producing promising and visually plausible results.
1 Introduction
@@
Given the rich semantics and complex inter-semantic relationships in the real-world scenes in Fig. 1(f), their available descriptions, shown in Fig. 1(a), are usually inadequately informative, thus leading to the semantic richness gap between the accordingly generated images and the real-world ones, as shown in Fig. 1(d) and (f). However, given the simple descriptions, our human minds can easily hallucinate a semantically abundant image, filling the semantic richness gap by first understanding the given description, then enriching it with its related semantics and inter-semantic relationships, and finally hallucinating the accordingly generated image.
The floor is under the sink. The floor has tile. The wall has a frame. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A tree and a street. A bush is next to the sidewalk. Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Input Description (b) Input Scene Graph (c) Enriched Scene Graph (d) Simple Generated (e) Enriched Generated (f) Real-World
Figure 1: Initially, the input textual description (a) is represented as a scene graph (b). Subsequently, we enrich the scene description (b) iteratively by appending additional objects and their relationships with existing scene elements. The enriching content (c) should preserve the same essential scene characteristics as the original input (b). This enrichment is executed utilizing our proposed end-to-end adversarial graph convolutional framework. Both the input (b) and enriched (c) scene graphs are then employed to synthesize simple (d) and enriched (e) images, respectively. In comparison to the simple image (d), the enriched one (e) not only reflects the essence of the initial input description (b) but also integrates more relevant intricate details akin to those present in real-world images (f) found in the Visual Genome dataset (Krishna et al., 2017).
@@
We term this task, barely investigated in conventional research, as generated contents enrichment (GCE). Although the human brain can fulfill this task effortlessly, existing artificial intelligence content generation methods (Karras et al., 2020; Vahdat & Kautz, 2020; Chen et al., 2020; Saharia et al., 2022; Rombach et al., 2022; Esser et al., 2021; Yang et al., 2022c; Herzig et al., 2020) produce results with inadequate semantic richness. In contrast to our approach, these methods overlook explicit content enrichment by neglecting reasoning on relevant semantics and inter-semantic relationships. As shown in Fig. 1(d), the state-of-the-art (SOTA) generation approach (Rombach et al., 2022), without reasoning the related semantics of the given descriptions, generates the images with insufficient semantics compared with the real-world ones in Fig. 1(f). Indeed, existing content generation methods, to the best of our knowledge, primarily rely on implicit learning to generate visual contents beyond those explicitly described in the input. However, this approach overlooks the semantic richness and structural reasonableness, making these methods inadequate for addressing GCE.
@@
Although GCE is essential for filling the semantic richness gap between the short description and the expected image, it is challenging to learn to solve it. Firstly, the enriching content should be semantically relevant to the ones in the description and form appropriate inter-semantic relationships. Secondly, the enriched scene graph, composited by the description and its enriched semantics, should be structurally real, following the distribution of scene graphs from real-world images.
@@
Towards solving the mentioned challenges and achieving GCE, we propose an intriguing approach that, to some extent, mimics the human reasoning procedure for hallucinating the enriching content. Particularly, our introduced method is designed with an end-to-end trainable architecture consisting of three principal stages. As an initial step, we represent the input description as a scene graph, wherein each node resembles a semantic, i.e., an object, and each edge corresponds to an inter-object relationship. In the first stage, we enrich the scene graph iteratively. In each iteration, we append a single object, representing it as an unknown node, along with its inter-object relationships as unknown edges. These are then predicted by employing a Graph Convolutional Network (GCN), which aggregates information from known nodes and edges. As a result, the enriching contents are semantically aligned with the ones in the input description. Then, the enriched scene graph is fed into a pair of Scene Graph Discriminators to guarantee its structural reality. In the second stage, we synthesize the enriched image from the enriched scene description by adopting an image generator (Johnson et al., 2018; Rombach et al., 2022). The third stage simultaneously feeds the enriched image and the input description into the Visual Scene Characteriser and Image-Text Aligner, thus the generated image is visually real and preserves the information of the description.
@@
Therefore, our contribution is the introduction of a novel adversarial end-to-end trainable framework designed to address the proposed GCE task as the first attempt, to our best knowledge. The GCE task aims to bridge the semantic richness gap between human hallucination and SOTA image synthesizers. Traditional text-to-image generation approaches solely focus on implicit visual enrichments. However, we enhance the generation by incorporating both explicit textual and visual content enrichments, ensuring semantic and structural coherence with the given input description. Our approach guarantees that the final enriched image accurately reflects the key scene elements specified in the input. We evaluate our method using real-world images and richly annotated scene descriptions from the Visual Genome dataset (Krishna et al., 2017), demonstrating promising quantitative and qualitative results.
2 Related Work
2.1 Realistic Image Generation
@@
Generative Adversarial Networks (GANs) (Goodfellow et al., 2020; Radford et al., 2015) jointly train a generator and discriminator to distinguish real from fake synthesized images. GANs produce images of high quality that are, in some cases, almost impossible for humans to discriminate from real images (Brock et al., 2018; Karras et al., 2020; Mirza & Osindero, 2014; Odena et al., 2017). However, their training procedure is not as stable as the other methods, and they may not model all parts of the data distribution (Metz et al., 2016). The first attempts for GAN-based text-to-image synthesis (Reed et al., 2016a; b) are followed by stacking GANs (Zhang et al., 2018) and employing the attention mechanism (Xu et al., 2018; Bahdanau et al., 2014; Vaswani et al., 2017). Variational Auto Encoder (VAE) (Kingma & Welling, 2013) consists of an encoder to a latent space and a decoder to return to the image space. VAE is also employed for image synthesis (Vahdat & Kautz, 2020; Huang et al., 2018), generating high-resolution data but with lower quality than their GAN counterpart. Alternatively, autoregressive approaches (Van Den Oord et al., 2016; Van den Oord et al., 2016; Chen et al., 2020; Child et al., 2019; Yu et al., 2022) synthesize pixels conditioned on an encoded version of the input description that could depend on the previously generated pixels with computationally expensive training and sequential sampling process. Although SOTA diffusion-based (Sohl-Dickstein et al., 2015) models (Dhariwal & Nichol, 2021; Ho et al., 2020; Song et al., 2020; Saharia et al., 2022) synthesize images of astounding quality, these architectures necessitate elevated computational costs during training and a time-consuming inference. Stable Diffusion (Rombach et al., 2022) is one of the first open-source methods that attracted so much attention among artists as it applies the diffusion process to low-dimensional image embeddings. On the other hand, there has been some research on a hybrid two-stage method like VQ-VAEs (Razavi et al., 2019; Yan et al., 2021) and VQGANs (Esser et al., 2021; Yu et al., 2021) which aims to benefit from a mixed version of the mentioned methods.
2.2 Scene Graphs
@@
Scene graphs proposed in Johnson et al. (2015a) and further investigated in Tripathi et al. (2021); Rosinol et al. (2021); Zhang et al. (2019) are a structured version of a scene representation containing information about objects, their relationships, and often their attributes. Previously, understanding a scene was limited to detecting and recognizing the objects in an image. However, scene graphs aim to obtain a higher level of scene understanding (Chang et al., 2021) and are employed to improve action recognition (Aksoy et al., 2010), image captioning (Aditya et al., 2015), and image retrieval (Johnson et al., 2015b). Scene graphs may be predicted from images (Lu et al., 2016; Xu et al., 2017; Yang et al., 2018; Tang et al., 2020; Yang et al., 2022b; Li et al., 2018) or serve as inputs to generate synthesized images (Johnson et al., 2018; Yang et al., 2022c; Herzig et al., 2020; Li et al., 2019; Ashual & Wolf, 2019).
2.3 Molecular Generation
Molecular generative models (Simonovsky & Komodakis, 2018; You et al., 2018; Shi et al., 2020; Luo et al., 2021; De Cao & Kipf, 2018; Goyal et al., 2020; Grover et al., 2019; Jin et al., 2018; Samanta et al., 2020) are designed specifically for molecular graphs, with a focus on applications in drug discovery. These models aim to generate molecular graphs that adhere to chemically valid structures while satisfying specific chemical property constraints. However, they are not directly applicable to scene graphs due to significant differences in structure. Molecular graphs are extremely smaller and less diverse, featuring fewer node types and relationships compared to scene graphs. Additionally, molecular generative models often assume that the input graph is connected, a property that does not necessarily hold for scene graphs (Agarwal et al., 2023).
2.4 Semantic Expansion
Our work introduces and tackles the task of GCE, which aims to enhance the quality of image synthesis. While prior research has explored various forms of semantic expansion, their goals, input-output structures, and evaluation criteria differ significantly from ours.
@@ 2.4 Semantic Expansion
The task of visual comprehension is addressed by generating an enriched and expressive scene graph as output from a given input image (Khan et al., 2023; 2022). These approaches incorporate common-sense knowledge derived from the input image to enhance the scene graph. Additionally, Yang et al. (2022a) focuses on the task of image outpainting, where an input image is extrapolated beyond its original boundaries by constructing and extending scene graphs based on the image. While these works are centered on generating scene graphs from input images, our GCE approach takes a different route by transforming a scene description into an enriched description and generating the corresponding enriched image.
GEMS (Agarwal et al., 2023) expands a seed scene graph for the purpose of image retrieval, iteratively generating novel scenes in the language domain using recurrent neural networks. In contrast, our approach enriches a scene graph entirely within the graph domain, with a clear focus on producing a visually enriched and coherent image. Unlike GEMS, we ensure the enriched scene graph incorporates objects and relationships that are specifically selected to enhance the quality and preserve the coherence of the synthesized image. This distinction is essential as we consider the impact of the enriching elements on the final generated image. In addition, both GCE and GEMS respect object co-occurrence patterns observed in the training data and are capable of generating diverse scene enrichments.
3 Method
@@
Figure 2: High-Level Overview of our end-to-end Generated Contents Enrichment framework during the training phase. In Stage 1, the input scene graph (SG) is fed to the Scene Graph Enricher ${G_{sg}}$ to produce an enriched SG. Besides, a pair of local and global discriminators in the SG Critic $D_{sg}$ are jointly trained to differentiate between original and enriched SGs. These discriminators aid the enrichment process in constructing realistic, structurally coherent, and semantically meaningful scenes. In Stage 2, the Image Synthesizer $G_{im}$ leverages the resultant enriched SG to generate an image. In Stage 3, essential visual and textual scene characteristics are extracted in the Visual Scene Characterizer $S_{cf}$ and the Image-Text Aligner $M_{im\_sg}$. These two components ensure that the enriched image appropriately reflects the original description’s inherent characteristics.
@@
Fig. 2 illustrates a high-level representation of our end-to-end model during the training phase. The model accepts a scene graph (SG) $g$ as input to simultaneously enrich the content in both textual and visual domains. Our proposed approach for GCE consists of three stages performed iteratively.
@@ 3.1 Scene Graphs
3.2 Graph Convolutional Network (GCN)
Our employed GCN is inspired by Johnson et al. (2018) and produces an output graph that maintains a structure analogous to the input graph. Each GCN is composed of multiple graph convolution (GConv) layers utilized to aggregate the data from nodes and edges, enabling the propagation of information throughout the graph. In this framework, the $l^{th}$ GConv layer, denoted as $GConv^{(l)}$ performs the message-passing as:
@@ 3.2 Graph Convolutional Network (GCN)
In the context of the $GConv^{(l)}$’s input graph, $V_{O}^{(l)}$ represents the feature vectors for all nodes, and $V_{R}^{(l)}$ denotes the feature vectors for all predicates. Additionally, the corresponding output vectors are characterized as $V_{O}^{(l+1)}$ and $V_{R}^{(l+1)}$. To initialize the process, $V_{O}^{(0)}$ and $V_{R}^{(0)}$ serve as the initial vector values for the objects and predicates within the GCN input graph.
Like CNNs, our approach employs weight sharing, enabling the GCN to accommodate graphs with an arbitrary number of vertices and edges. Additionally, because the computations for each node are independent, parallelization is achievable, which enhances overall computational efficiency.
The fundamental building block of our GCN is the GConv layer. For every object $o_{i}\in O$ and edge $(o_{i},r,o_{j})\in E$, we concatenate the three corresponding sets of feature vectors, which consist of $(v_{i},v_{r},v_{j})$ representing the subject, predicate, and object vectors, respectively. These concatenated feature sets are then passed through the GConv layer for further processing, initiating with an MLP represented as $f_{g}$, where $(\bar{v}_{i},\bar{v}_{r},\bar{v}_{j})=f_{g}(v_{i},v_{r},v_{j})$. This operation computes the set of three intermediate subject, predicate, and object vectors $(\bar{v}_{i},\bar{v}_{r},\bar{v}_{j})$. Subsequently, these intermediate vectors are aggregated to form $V_{O}^{(l)}$ and $V_{R}^{(l)}$. Node features $V_{O}^{(l)}$ include the new feature vectors $v^{\prime}_{i}$ for the output nodes and edge features $V_{R}^{(l)}$ consist of the new feature vectors $v^{\prime}_{r}$ for the output predicates:
$v^{\prime}_{r}=f_{r}(v_{r},\bar{v}_{r}),\;\;\;\;C_{i}^{s}=\{\bar{v}_{i}|(\bar{v}_{i},\bar{v}_{r},\bar{v}_{j})=f_{g}(v_{i},v_{r},v_{j})\text{ and }(o_{i},r,o_{j})\in E\},$ (2)
$\begin{split}v^{\prime}_{i}=f_{o}(v_{i},\frac{1}{|C_{i}^{s}|+|C_{i}^{o}|}\sum_{\bar{v}_{i}\in C_{i}^{s}\cup C_{i}^{o}}\bar{v}_{i}),\;\;\;\;C_{i}^{o}=\{\bar{v}_{i}|(\bar{v}_{j},\bar{v}_{r},\bar{v}_{i})=f_{g}(v_{j},v_{r},v_{i})\text{ and }(o_{j},r,o_{i})\in E\}.\end{split}$ (3)
@@
For more comprehensive information regarding the structure of our GCN and GConv, including detailed illustrations and further explanations, please refer to the supplementary material.
3.3 Stage 1: Generative Adversarial SG Enrichment
3.3.1 SG Enricher
The Scene Graph Enricher, serving as our SG generator and featured in Fig. 2, encompasses three distinct steps, each tailored to address specific subproblems associated with appending a new object, edges, and predicates. These steps leverage the proposed GCNs along with an Enriching Edge Detector. Here is an overview of the process: